OpenStudy (anonymous):

If A=area of a circle with radius r and the circle expands as time passes, find dA/dt in terms of dx/dt? I'm not sure how to do this

OpenStudy (anonymous):

me neither, since there is no \(x\) in \(A=\pi r^2\)

OpenStudy (anonymous):

Related Rates

OpenStudy (anonymous):

do you mean in terms of dr/dt?

OpenStudy (anonymous):

She must mean mean dr/dt...

OpenStudy (anonymous):

oh, wow. Sorry! I meant in terms of dr/dt!

OpenStudy (anonymous):

have you heard of "related rates" before?

OpenStudy (anonymous):

Actually, you don't even need to get into that--just use the chain rule.

OpenStudy (anonymous):

You know it?

OpenStudy (anonymous):

Yes, I got 2pir(dr/dt)

OpenStudy (anonymous):

that=dA/dt

OpenStudy (anonymous):

Yeah, that's great! So what's your question?

OpenStudy (anonymous):

My actual question was "Suppose oil spills from a ruptured tanker and spreads in a circular pattern. If the radius of the oil spill increases at a constant rate of 1 m/s, how fast is the area of the spill increasing when the radius is 30 m?"

OpenStudy (anonymous):

Ok, now you know dr/dt = 1 m/s. And also that r=30m. So just plug those into the formula you got for dA/dt

OpenStudy (anonymous):

I got 60pim

OpenStudy (anonymous):

*m/s

OpenStudy (anonymous):

Yeah, that sounds good to me...

OpenStudy (anonymous):

Sorry, thanks so much for your help!

OpenStudy (anonymous):

No problem-- You did all the work!